list the first five terms of the arithmetic sequence.
A1= -3 ,An+1=An+6

Answer by stanbon(75887)

list the first five terms of the arithmetic sequence.
A1= -3 ,
An+1 = An+6
----
Ans:: -3,-3+6=3,3+6=9,15,21,27
-----------
Cheers,
Stan H.